what has a greater distance ? a Square with a length of 9ft or a radius with a distance of 5ft ?
Answer by stanbon(75887)

what has a greater distance ? a Square with a length of 9ft or a radius with a distance of 5ft ?
----
Perimeter of the square = 4*9 = 36 ft
---
Perimeter of the circle = 2(pi)r = 10*pi = 31.4.. ft
========================================================
Cheers,
Stan H.